Using Iror (high cost over low cost), Show your work finding NPW and make recommendation Also do CFDs for the options and another one for the Delta between them (High cost - Low cost) Both options have useful life of 3 years

A B
First cost $100 $140
Cost at year 3 $100 $0

Take MARR = 30%
Answer #1

CFO for A MARR = 30% lo -10 CFO ar B w -10 NPW PW of costs = Pw of Benefits - loo for = [ 100 + 1 1.3 NOW А = -10ot 45.52 = 8Delta cash flow diagram + $100 o - $40 o = High cost - low cost -140- (-100) = -40 o fint cost o cast year 3 = o-(-1) = 100 (